OpenStudy (anonymous):

The length of a rectangle is three times its width. If the perimeter is at most 112 centimeters, what is the greatest possible value for the width? Solve the problem.

OpenStudy (anonymous):

A.) 42 cm B.)22.4 cm C.) 14 cm D.) 28 cm

OpenStudy (anonymous):

14 cm

OpenStudy (anonymous):

l = 3b 2(l+b) = 112 2(3b+b) = 112 2(4b)=112 4b=56 b=14

OpenStudy (ash2326):

l=3b l=length b=breadth perimeter at most is 112 perimeter=2(l+b) 2(l+b)=112 l+b=56 l=3b 3b+b=56 4b=14 b=14 that's the max value of width

OpenStudy (anonymous):

4b = 56 @ash